# LLaVA-OneVision-1.5: Fully Open Framework for Democratized Multimodal Training

### v1

@@ 5.1 Overall Performance
We use LMMs-Eval Zhang et al. (2024) with the default prompt to evaluate the performance of LLaVA-OneVision-1.5 across multiple benchmarks in four categories of downstream tasks:(1) General Visual Question Answering (VQA): MMStar (Chen et al., 2024a), MMEBench series (Fu et al., 2023), MME-RealWorld series (Zhang et al., 2025), SeedBench (Li et al., 2024b), Seed-Bench-2-Plus (Li et al., 2024a), CV-Bench (Tong et al., 2024), and RealWorldQA (Corp., 2024). (2) Multimodal Reasoning: MathVista (Lu et al., 2024), WeMath (Qiao et al., 2025), MathVision (Wang et al., 2024a), MMMU (Yue et al., 2024), and MMMU-Pro series (Yue et al., 2025).(3) OCR & Chart Understanding: ChartQA (Masry et al., 2022), CharXiv (Wang et al., 2024d), DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2024c), AI2D (Kembhavi et al., 2016), and InfoVQA (Mathew et al., 2022).(4) Others: PixmoCount (Deitke et al., 2025), CountBench (Paiss et al., 2023), VL-RewardBench (Li et al., 2025b), and V∗ (Wu and Xie, 2024). As shown in Tab. 1, LLaVA-OneVision-1.5-8B surpasses Qwen2.5-VL-7B on 18 of 27 benchmarks and LLaVA-OneVision-1.5-4B surpasses Qwen2.5-VL-3B on 27 of 27 benchmarks.
@@ 5.6.1 Comparison of Different Vision Encoders
In Tab. 2, we evaluate various vision encoders, CLIP (Radford et al., 2021), MLCD (An et al., 2024), AIMv2 (Fini et al., 2025), DFN (Fang et al., 2023), SigLIP (Zhai et al., 2023), SigLIPv2 (Tschannen et al., 2025), and RICE-ViT (Xie et al., 2025), within the LLaVA-NeXT framework. At a resolution of 336 pixels, RICE-ViT surpasses CLIP across all benchmarks, achieving significant gains in InfoVQA (+6.3%) and OCRBench (+5.0%). It also demonstrates significant improvements in document understanding compared to AIMv2. At 378 pixels, RICE-ViT outperforms computationally intensive models such as SigLIPv2 in 9 of 14 benchmarks, notably in InfoVQA (+4.4%), DocVQA (+3.5%), and ChartQA (+4.9%). These results position RICE-ViT as a leading vision encoder, enhancing OCR capabilities and providing robust visual understanding, crucial for applications requiring advanced document analysis and visual reasoning. In addition, we further compare the performance of RICE-ViT with that of Qwen-ViT after incorporating LMM training, where RICE-ViT is derived from LLaVA-OneVision-1.5-3B and Qwen-ViT is derived from Qwen2.5-VL-7B. In the areas of OCR & Document Understanding and General Vision Understanding, RICE-ViT demonstrates average performance improvements of 1.9% and 0.9% compared to Qwen-ViT.
@@ 5.6.4 Instruction Data Quality and Scaling
To enhance performance across diverse VQA tasks, we compile 124 types of instruction data (LLaVA-OneVision-1.5-Inst-Data) for SFT training. We further scale the model capabilities by deduplicating and merging the recently proposed FineVision dataset (HuggingFaceM4, 2025), resulting in the Merged46M SFT dataset. To maintain consistent training steps, we double the batch size for Merged46M due to its larger scale. Fig. 7 shows performance comparisons on 16 benchmarks during SFT using three datasets: LLaVA-OneVision-1.5-Inst-Data, FineVision, and Merged46M. LLaVA-OneVision-1.5-Inst-Data achieves performance comparable to FineVision, while the Merged46M dataset delivers the best results across nearly all benchmarks.

### v3

@@ 6.1 Overall Performance
We use LMMs-Eval lmms_eval with the default prompt to evaluate the performance of LLaVA-OneVision-1.5 across multiple benchmarks in four categories of downstream tasks:(1) General Visual Question Answering (VQA): MMStar (mmstar), MMEBench series (mme), MME-RealWorld series (mmerealworld), SeedBench (seedbench), Seed-Bench-2-Plus (seedbench2), CV-Bench (tong2024cambrian), and RealWorldQA (realworldqa). (2) Multimodal Reasoning: MathVista (mathvista), WeMath (wemath), MathVision (mathvision), MMMU (mmmu), and MMMU-Pro series (mmmupro).(3) OCR & Chart Understanding: ChartQA (chartqa), CharXiv (charxiv), DocVQA (docvqa), OCRBench (ocrbench), AI2D (ai2d), and InfoVQA (infovqa).(4) Others: PixmoCount (molmo), CountBench (paiss2023teaching), VL-RewardBench (li2025vl), and V∗ (wu2024v). As shown in Tab. 1, LLaVA-OneVision-1.5-8B surpasses Qwen2.5-VL-7B on 18 of 27 benchmarks and LLaVA-OneVision-1.5-4B surpasses Qwen2.5-VL-3B on 27 of 27 benchmarks.
@@ 6.5 Others
6.6 Performance of RL Post-training
To validate the effectiveness of our lightweight RL post-training, we compare the RL-enhanced models (LLaVA-OV-1.5 RL-8B) against their supervised baselines and the strong competitor Qwen2.5-VL-7B. We report results for both standard "fast" inference and "thinking" mode (where the model generates reasoning chains) as shown in Tab. 1.
Core Capability Enhancement. As presented in Tab. 1, RL post-training yields consistent gains across major benchmarks. (1) Multimodal Reasoning: The most substantial improvements are observed in multimodal reasoning tasks. In "thinking" mode, our model achieves dramatic gains on WeMath (+7.9), MathVision (+8.8), and MMMU-Pro (+10.5), demonstrating that the RL-induced chain-of-thought capability effectively unlocks deeper problem-solving skills. (2) General VQA & OCR: On standard benchmarks like MMBench and DocVQA, RL maintains or slightly improves the already strong SFT performance, ensuring no regression in general capabilities while specializing in hard reasoning.
Extended Capability Analysis. Beyond the core benchmarks, we utilize Figure 5 to analyze the impact of RL on specific vertical capabilities not fully covered in Tab. 1. (1) Spatial Reasoning & Grounding: As shown in Figure 5, RL significantly enhances fine-grained perception. The RL model (Fast mode) consistently outperforms the SFT baseline on spatial tasks like SAT and Ref-L4. Interestingly, while "thinking" mode aids reasoning, it sometimes yields lower scores than "fast" mode on strictly perceptual metrics (e.g., Ref-L4 IoU), suggesting that verbose generation may occasionally interfere with precise coordinate regression. (2) Coding: In the coding domain (WebCode, UniSVG), Figure 5 shows that our RL models achieve consistent gains. The "thinking" mode proves particularly effective here, achieving the highest scores on Design2Code and UniSVG, indicating that chain-of-thought reasoning is beneficial for structural code generation.
These results confirm that our rule-based RL framework effectively enhances both the reasoning depth (via thinking chains) and perceptual precision (via direct feedback) of LLaVA-OneVision-1.5, positioning it as a state-of-the-art open model in the 8B parameter class.
@@ 6.7.1 Comparison of Different Vision Encoders
In Tab. 2, we evaluate various vision encoders, CLIP (clip), MLCD (mlcd), AIMv2 (fini2025multimodal), DFN (dfn), SigLIP (siglip), SigLIPv2 (siglipv2), and RICE-ViT (rice), within the LLaVA-NeXT framework. At a resolution of 336 pixels, RICE-ViT surpasses CLIP across all benchmarks, achieving significant gains in InfoVQA (+6.3%) and OCRBench (+5.0%). It also demonstrates significant improvements in document understanding compared to AIMv2. At 378 pixels, RICE-ViT outperforms computationally intensive models such as SigLIPv2 in 9 of 14 benchmarks, notably in InfoVQA (+4.4%), DocVQA (+3.5%), and ChartQA (+4.9%). These results position RICE-ViT as a leading vision encoder, enhancing OCR capabilities and providing robust visual understanding, crucial for applications requiring advanced document analysis and visual reasoning. In addition, we further compare the performance of RICE-ViT with that of Qwen-ViT after incorporating LMM training, where RICE-ViT is derived from LLaVA-OneVision-1.5-3B and Qwen-ViT is derived from Qwen2.5-VL-7B. In the areas of OCR & Document Understanding and General Vision Understanding, RICE-ViT demonstrates average performance improvements of 1.9% and 0.9% compared to Qwen-ViT.
@@ 6.7.4 Instruction Data Quality and Scaling
To enhance performance across diverse VQA tasks, we compile 124 types of instruction data (LLaVA-OneVision-1.5-Inst-Data) for SFT training. We further scale the model capabilities by deduplicating and merging the recently proposed FineVision dataset (finevision2025), resulting in the Merged46M SFT dataset. To maintain consistent training steps, we double the batch size for Merged46M due to its larger scale. Fig. 8 shows performance comparisons on 16 benchmarks during SFT using three datasets: LLaVA-OneVision-1.5-Inst-Data, FineVision, and Merged46M. LLaVA-OneVision-1.5-Inst-Data achieves performance comparable to FineVision, while the Merged46M dataset delivers the best results across nearly all benchmarks.